Passing the ball between two players during a soccer game is a critical skill for the success of a team. A random sample of passes made by the English and German teams in the 2010 World Cup was drawn, and the number of successful passes in each sample was counted. See data below.

Germany England
# of successful passes 69 71
Total in the sample 76 91

Consider the soccer problem above. The German coach wants to change the conclusion of this study to be a conclusion in his favor. What could he do?

Decrease both sample sizes

Change the level of significance

There is nothing he can do to change the conclusion

Use paired observations
Answer #1

Ans:

sample proportion 1=69/76=0.9079

sample proportion 2=71/91=0.7802

Test statistic:

z=(0.9079-0.7802)/SQRT(0.8383*(1-0.8383)*((1/76)+(1/91)))

z=2.232

p-value=P(z>2.23)=0.0128

Correct option is:

Change the level of significance

We can increase the level of significance for getting significance level,so that we can reject the null hypothesis.